OpenStudy (anonymous):

7f+g=5 solve for f **show work**

OpenStudy (anonymous):

here are my steps 1) subtract \(g\) from both sides 2) divide both sides by \(7\)

OpenStudy (anonymous):

7f+g=5 or, 7f=5-g or, f=(5-g)/7 so, f=(5-g)/7

OpenStudy (anonymous):

i dont knw if its corect or nt!